anna weighs 3kg more than her sister and their total weight is 105kg,what is the weigth of each girl
Answer by alicealc(293)

anna's weight = a
her sister's weight = s

a = 3 + s

a + s = 105
3 + s + s = 105
3 + 2s = 105
2s = 105 - 3
2s = 102
s = 102/2
s = 51
a = 3 + 51 = 54

so, Anna's weight = 54 kg, her sister's weight = 51 kg